OpenStudy (anonymous):

Prove Trig Equation: PLEASE HELP ME!!! -tan^2x + sec^2x = 1

OpenStudy (goformit100):

-tan^2x + sec^2x = 1 or, sec^2x -tan^2x = 1 or, (h/b)^2 - (p/b)^2 =1 or, h^2/b^2 - p^2/b^2 = 1 or, (h^2 - p^2)/b^2 = 1 or, b^2/b^2 =1...................proved

OpenStudy (anonymous):

Wait so 'or' means all the different ways to prove? @goformit100

OpenStudy (goformit100):

no it's the substitute for 'emplies that'

OpenStudy (anonymous):

oh oops okay so thats how you would prove it?

OpenStudy (goformit100):

Yes Ma'am.

OpenStudy (anonymous):

im confused on what the h, b, and p represents? sorry Im just trying to understand.

OpenStudy (goformit100):

h means hypotenuse of the right angled triangle b is the base and p is the perpendicular

OpenStudy (anonymous):

So you could use a triangle to solve this or is it just a saying kind of thing?

OpenStudy (goformit100):